# Control of photoionization by resonant phase-locked pulse pairs

### v1

@@ IV.1 Population dynamics
First, we investigate phase-dependent population dynamics of helium using RTDCIS for the two-pulse case during resonant photoionization (1+1). Linearly polarized pulses are described by the vector potential
$A(t)=A_{0}\sin[\omega(t-t_{0})-\varphi]\cos\left[\frac{\pi(t-t_{0})}{\tau}\right]^{2}\Theta,$ (11)
where $A_{0}$ is the amplitude, $\omega$ is the carrier frequency, $\varphi$ is the carrier envelope phase (CEP), $t_{0}$ is the center of the pulse, and $\tau$ is the foot-to-foot pulse duration, where $\Theta=1$ for $|t-t_{0}|<\tau/2$ and 0 otherwise. Both pulses are taken to have the same parameters with the electric field $E(t)=-\dot{A}(t)$ corresponding to a pulse area of $\theta\approx 2.5\pi$. The probe pulse is delayed by a fixed amount $t_{0}>\tau$ with a variable CEP $\varphi\in[0,2\pi]$, while the pump pulse is fixed to give a zero reference at $t_{0}^{\mathrm{pump}}=0$ with $\varphi^{\mathrm{pump}}=0$. In this configuration, the field phase of the probe field becomes
$\phi=\omega t_{0}+\varphi,$ (12)
which is expected to capture the electron dynamics on dressed states during the probe pulse for the values $(2N+1)\frac{\pi}{2}$, as explained in Sec. III. For time-dependent simulations, a complex absorbing potential (CAP) was used instead of ECS to enforce outgoing boundary conditions (since ECS can only be used reliably for time-dependent simulations in the velocity gauge, but there TDCIS suffers from unphysical many-body energy shifts [68]). The form of the CAP used was
@@
RTDCIS includes all states with single excitations. We find that the triplet state $1s2p\,^{3}P_{1}$ is not populated, which is to be expected due to the absent coupling to the singlet ground state due to symmetry restrictions in the present scenario. Hence, there are no qualitative differences between the relativistic and non-relativistic results for helium, at the level of RTDCIS and TDCIS, and it is not necessary to make a detailed comparison between the two methods in this work.
@@ IV.2 Photoelectron dynamics
In this subsection, we present numerical results for energy-resolved photoionization of helium atoms for the phase-locked two-pulse, resonant two-photon ionization (1+1) process, shown in Figure 5. As will be explained below, our results show that control of the dressed state population by $\varphi$ manifests itself not only in the energy-resolved photoelectron signal, but also in the total amount of ionization.
The simulations were performed with TDCIS and t-SURFF in the length gauge using both linear and circular polarization. Both pulses have the same duration, but the intensity is increased to the stabilization criterion for circular polarization for helium at $10^{14}$ W/cm2, see Fig. 2 (d). The resulting pulse areas are approximately $3.5\pi$, i.e., the required $\pi/2$-pulse type, leading to nearly equal populations of the two essential atomic states by the pump pulse. Photoelectron spectra were calculated using the Zonte t-SURFF implementation [74]. The surface flux was evaluated at a radius of $r=92$ a.u. with a maximum single particle angular momentum of $\ell_{\textrm{max}}=7$.
@@ IV.2.1 Photoelectron spectra
In our approach, the second pulse introduces additional structures in the photoelectron spectrum. Ramsey-like fringes appear with a spacing that depends on the inverse delay, $t_{0}>\tau$, between pump and probe pulses, similar to that in Ref. [11]. This yields an interference signal for photoionization that depends on both the pump and the probe process. The contrast of these fringes gives an indication of the relative ionization probabilities from the two pulses. At $\varphi=4.40$, the fringes show less contrast, confirming the reduction in ionization during the probe pulse by targeting the $\left|+\right\rangle$ state. Lineouts from Fig. 5 (a) are presented in Fig. 5 (b), showing a clear transition from high contrast ($\varphi=1.26$) to strongly reduced contrast ($\varphi=4.4$). Additionally, the CEP value $\varphi=0$ is found to display nearly maximal contrast. This can only be attained when the two photoelectron sources are equally strong, which in our case implies that the atom must Rabi cycle in the same way driven by both the pump and the probe field. Analysis of this particular CEP value, $\varphi=0$, using Eq. (12), reveals that it corresponds to a field phase of $\phi=1.13\pi\approx\pi$. The interpretation is then clear: the probe pulse rewinds the atomic dynamics back to the ground state making the ionization process inversion symmetric with that of the pump pulse. Recently, such time symmetries have been exploited for control of quantum entanglement in photoionization and in entanglement transfer from ions to spontaneous photons by Stenquist et al. in Refs. [75, 76].
We mention that, prior to our work, the case of circular pulses with opposite helicity between the pump and probe has been shown to transform the Ramsey fringes such that they are reduced (or eliminated) without angular resolution. This is because the photoelectron momentum distribution (PMD) would exhibit spiral patterns in the plane of the polarization of the light [77]. Photoelectron spectra along a given azimuth, however, would still show readily apparent fringes in the case that the second pulse ionizes, a pattern against which stabilized results could be contrasted just like for co-rotating circularly polarized setup. Also, in addition to the AT doublet (or single peak in our case), a third, central (uncoupled) photoelectron peak can appear in the case of two pulses with opposite helicity as pointed out by Deng et al. in 2025 [78], these effects are not present in our scheme because co-rotating pulses are employed.
@@ IV.2.2 Ionization ratios
Finally, the linearly polarized case in Fig. 5 (c) with $10^{14}$ W/cm2 has the same pulse area, $\theta\approx 3.5\pi$, as the circular case with the same intensity. This makes it a good test case for the theory of stabilization. As should be expected from Fig. 2 (c) and (d), the linearly polarized case shows a reduced amount of control over the ionization, which we attribute to the multiple continua (s- and d-waves) that are coupled from the essential states. The fact that the difference is not larger between the two cases is probably due to our choice of a smooth $\cos^{2}$-envelope function in Eq. (11), while a stronger stabilization would likely arise with a more rectangular pulse. A systematic study of such envelope effects could be done using the super-Gaussian sequence, as was recently done for ultrafast absorption and resonance fluorescence at XUV wavelengths [79]), but such a detailed analysis remains beyond the scope of the present work.
@@ IV.2.3 Ionization over time
Finally, Fig. 5 (d) shows the time-resolved photoelectron flux generated by the pump-probe sequence for the same values of $\varphi$ included in panel (b). It is observed that the photoionization from the probe pulse can be controlled with suppression or enhancement of the total yield. Field phases separated by $\pi$ give maximal enhancement or suppression. This means that the energy resolution of the photoelectron is not required to observe the proposed non-adiabatic stabilization phenomenon. In practice, this is important because the Ramsey-like fringes are likely to be washed out by macroscopic averaging effects in real experiments.
@@ Acknowledgements.
Appendix A Gauge invariance

### v2

@@ IV.1 Population dynamics
First, we investigate phase-dependent population dynamics of helium using RTDCIS for the two-pulse case during resonant photoionization (1+1). Linearly polarized pulses, as defined in Sec. III.1, are used. Both pulses are taken to have a pulse area of $\theta\approx 2.5\pi$. For time-dependent simulations, a complex absorbing potential (CAP) was used instead of ECS to enforce outgoing boundary conditions (since ECS can only be used reliably for time-dependent simulations in the velocity gauge, but there TDCIS suffers from unphysical many-body energy shifts [7]). The form of the CAP used was
@@
RTDCIS includes all states with single excitations. We find that the triplet state $1s2p\,^{3}P_{1}$ is not populated, which is expected due to the absence of coupling to the singlet ground state because of symmetry restrictions in the present scenario. Hence, there are no qualitative differences between the relativistic and non-relativistic results for helium, at the level of RTDCIS and TDCIS, and it is not necessary to make a detailed comparison between the two methods in this work.
@@ IV.2 Photoelectron dynamics
In this subsection, we present numerical results for energy-resolved photoionization of helium atoms for the phase-locked two-pulse, resonant two-photon ionization (1+1) process, as shown in Figure 5. As will be explained below, our results show that control of the dressed state population by $\varphi$ manifests itself not only in the energy-resolved photoelectron signal, but also in the total amount of ionization.
The simulations were performed with TDCIS and t-SURFF in the length gauge using both linear and circular polarization. Both pulses have the same duration, but the intensity is increased to meet the stabilization criterion for circular polarization for helium at $10^{14}$ W/cm2, see Fig. 2 (d). The resulting pulse areas are approximately $3.5\pi$, i.e., the required $\pi/2$-pulse type, leading to nearly equal populations of the two essential atomic states by the pump pulse. Photoelectron spectra were calculated using the Zonte t-SURFF implementation [8]. The surface flux was evaluated at a radius of $r=92$ a.u. with a maximum single particle angular momentum of $\ell_{\textrm{max}}=7$.
@@ IV.2.1 Photoelectron spectra
In our approach, the second pulse introduces additional structures in the photoelectron spectrum. Ramsey-like fringes appear with a spacing that depends on the inverse delay, $t_{0}>\tau$, between pump and probe pulses, similar to that in Ref. [74]. This yields an interference signal for photoionization that depends on both the pump and the probe process. The contrast of these fringes gives an indication of the relative ionization probabilities from the two pulses. At $\varphi=4.40$, the fringes show less contrast, confirming the reduction in ionization during the probe pulse by targeting the $\left|+\right\rangle$ state. Lineouts from Fig. 5 (a) are presented in Fig. 5 (b), showing a clear transition from high contrast ($\varphi=1.26$) to strongly reduced contrast ($\varphi=4.4$). Additionally, the CEP value $\varphi=0$ is found to display nearly maximal contrast. This can only be attained when the two photoelectron sources are equally strong, which, in our case, implies that the atom must Rabi cycle in the same way driven by both the pump and the probe field. Analysis of this particular CEP value, $\varphi=0$, using Eq. (6), reveals that it corresponds to a field phase of $\phi=1.13\pi\approx\pi$. The interpretation is then clear: the probe pulse rewinds the atomic dynamics back to the ground state, making the ionization process inversion symmetric with that of the pump pulse. Recently, such time symmetries have been exploited for the control of quantum entanglement in photoionization and in entanglement transfer from ions to spontaneous photons by Stenquist et al. in Refs. [67, 66].
We mention that, prior to our work, the case of circular pulses with opposite helicity between the pump and probe has been shown to transform the Ramsey fringes such that they are reduced (or eliminated) without angular resolution. This is because the photoelectron momentum distribution (PMD) would exhibit spiral patterns in the plane of the polarization of the light [28]. Photoelectron spectra along a given azimuth, however, would still show readily apparent fringes in the case that the second pulse ionizes, a pattern against which stabilized results could be contrasted just like for a co-rotating circularly polarized setup. Also, in addition to the AT doublet (or single peak in our case), a third, central (uncoupled) photoelectron peak can appear in the case of two pulses with opposite helicity, as pointed out by Deng et al. in 2025 [16], these effects are not present in our scheme because co-rotating pulses are employed.
@@ IV.2.2 Ionization ratios
Finally, the linearly polarized case in Fig. 5 (c) with $10^{14}$ W/cm2 has the same pulse area, $\theta\approx 3.5\pi$, as the circular case with the same intensity. This makes it a good test case for the theory of stabilization. As should be expected from Fig. 2 (c) and (d), the linearly polarized case shows a reduced amount of control over ionization, which we attribute to the multiple continua (s- and d-waves) that are coupled from the essential states. The fact that the difference is not larger between the two cases is probably due to our choice of a smooth $\cos^{2}$-envelope function in Eq. (3), while a stronger stabilization would likely arise with a more rectangular pulse. A systematic study of such envelope effects could be conducted using the super-Gaussian sequence, as was recently done for ultrafast absorption and resonance fluorescence at XUV wavelengths [68], but such a detailed analysis remains beyond the scope of the present work.
@@ IV.2.3 Ionization over time
Finally, Fig. 5 (d) shows the time-resolved photoelectron flux generated by the pump-probe sequence for the same values of $\varphi$ included in panel (b). It is observed that photoionization from the probe pulse can be controlled with the suppression or enhancement of the total yield. Field phases separated by $\pi$ give maximal enhancement or suppression. This means that the energy resolution of the photoelectron is not required to observe the proposed non-adiabatic stabilization phenomenon. In practice, this is important because the Ramsey-like fringes are likely to be washed out by macroscopic averaging effects in real experiments.
@@ Acknowledgements.
Appendix A Gauge invariance and restricted calculations
@@
Appendix B Comparison of dipole moments between CIS and SAE
Consider the two orbitals $\phi_{1s}$ and $\phi_{2p}$. Within CIS and the spin-singlet approximation, we can represent the ground-state and a singly excited state by the following states
$\left|\Psi_{1s^{2}}\right\rangle$ $=\frac{1}{\sqrt{2}}\left|\phi_{1s},\phi_{1s}\right\rangle(\left|\uparrow\downarrow\right\rangle-\left|\uparrow\downarrow\right\rangle),$ (17)
$\left|\Psi_{1s2p}\right\rangle$ $=\frac{1}{2}(\left|\phi_{1s},\phi_{2p}\right\rangle+\left|\phi_{2p},\phi_{1s}\right\rangle)\left(\left|\uparrow\downarrow\right\rangle-\left|\uparrow\downarrow\right\rangle\right).$ (18)
The matrix element for the $z$-component of the dipole moment $z=z_{1}+z_{2}$, where $z_{i}$ denotes the dipole moment of electron $i$, is given by
$\begin{split}\left\langle\Psi_{1s2p}\right|z\left|\Psi_{1s^{2}}\right\rangle&=\left\langle\Psi_{1s2p}\right|z_{1}\left|\Psi_{1s^{2}}\right\rangle\\ &+\left\langle\Psi_{1s2p}\right|z_{2}\left|\Psi_{1s^{2}}\right\rangle.\end{split}$ (19)
Since $z$ is spin-independent, the spin part of the state simply contributes an overall factor of 2 to the matrix element. The individual terms then give
$\left\langle\Psi_{1s2p}\right|z_{1}\left|\Psi_{1s^{2}}\right\rangle$ $=\frac{1}{\sqrt{2}}\left\langle\phi_{2p}\right|z_{1}\left|\phi_{1s}\right\rangle,$ (20)
$\left\langle\Psi_{1s2p}\right|z_{2}\left|\Psi_{1s^{2}}\right\rangle$ $=\frac{1}{\sqrt{2}}\left\langle\phi_{2p}\right|z_{2}\left|\phi_{1s}\right\rangle,$ (21)
which leads to
$\left\langle\Psi_{1s2p}\right|z\left|\Psi_{1s^{2}}\right\rangle=\sqrt{2}\left\langle\phi_{2p}\right|z\left|\phi_{1s}\right\rangle,$ (22)
where the electron index has been dropped on the RHS. If we instead consider an SAE model where the orbitals are the same as in CIS, then the corresponding matrix element would give
$\left\langle\Psi_{2p}\right|z\left|\Psi_{1s}\right\rangle=\left\langle\phi_{2p}\right|z\left|\phi_{1s}\right\rangle.$ (23)
Comparing Eq. (22) and (23), we can see the factor of $\sqrt{2}$ that is discussed in the main text. If we consider a third orbital, say $\phi_{3d}$, then we find instead
$\left\langle\Psi_{1s3d}\right|z\left|\Psi_{1s2p}\right\rangle=\left\langle\phi_{3d}\right|z\left|\phi_{2p}\right\rangle$ (24)
$\left\langle\Psi_{3d}\right|z\left|\Psi_{2p}\right\rangle=\left\langle\phi_{3d}\right|z\left|\phi_{2p}\right\rangle,$ (25)
and there is no difference between the models.